# CPEG: Leveraging Consistency Policy with Consensus Guidance for Multi-agent Exploration

### v1

@@ IV-B Consensus Guidance
The encoder $z_{e}(\cdot;\alpha_{e})$, parameterized by $\aleph_{e}$, encodes the observation of agent $o_{a}$ into an embedding $z_{e,a}\in\mathbb{R}^{m}$, matching the length of the consensus representation $m$. In practice, we use historical observations as inputs during each agent’s execution to capture sufficient information about the global state.
Consensus Codebook
The process of inferring consensus for each input observation proceeds as follows. First, the embedding $z_{e,a}$ is mapped to a discrete index $k_{a}$ of the codebook $\mathcal{E}$:
@@
where $e_{j}\in\mathcal{E}$ is a codebook vector, and $K$ denotes the size of the codebook. Notably, the performance of consensus learning is relatively insensitive to the choice of $K$, allowing us to set the number of discrete consensus in codebook $K=5$ across all scenarios. Based on this discrete index $k_{a}$, the discrete consensus representation for the agent $a$ is derived as $\psi_{a}=e_{k_{a}}$. The loss function for the consensus inference process is defined as:
@@
where $\mathrm{sg}$ denotes the straight-through gradient operation [39], enabling back-propagation training of $\operatorname*{\arg\!\min}$ operations, and $\beta$ is a hyperparameter balancing the degree of alignment between the codebook and input embeddings. In practice, CPEG uses the exponential moving average (EMA) to update the codebook, instead of directly learning them as parameters. Specifically, EMA updates the codebook $e_{k_{a}}$ by replacing it with a weighted combination of its previous value $e_{k_{a}}$ and the corresponding vector $z_{e,a}$:
@@
To develop the ability to deduce consensus from local observations, we aggregate the consensus $\psi_{a}$ from all agents together and process them through a state decoder $z_{d}(\cdot;\alpha_{d})$, parameterized by $\alpha_{d}$, to reconstruct the global state $\hat{s}$. We train the observation encoder and state decoder by maximizing the log-likelihood of the reconstruction:
@@
which encourages the encoding and decoding processes of the consensus learner to capture the key information from partial observations for accurate consensus deduction. In addition, the reconstruction loss gradient is also passed to the encoder for training $\alpha_{e}$ by straight-through gradient estimation.
Mask Consensus Exploration
In this work, we employ two distinct exploration strategies: unguided exploration and consensus-guided exploration, aiming to balance exploration and exploitation to some degree. Unguided exploration allows agents to act according to their current policies without specific consensus guidance, emphasizing their self-directed exploration capabilities. However, its effectiveness is limited by the initial policy quality and the inherent complexities of MAS, neglecting potential cooperative advantages. Considering this, we introduce another exploration strategy, consensus-guided exploration designed to steer the cooperation efforts of agents. The consensus-guided exploration instructs agents to take actions based on the consistency policy guided by their assigned consensus. Through the guidance of shared consensus among different agents, consistency policies can generate cooperative exploration behaviors. Meanwhile, to incorporate consensus guidance, the consistency function $F_{\theta}(o,u,\tau)$ in Eq. 4 is modified to $F_{\theta}(o,u,\psi,\tau)$.
To implement the above concept and combine the two exploration mechanisms, we introduce a binary consensus mask $\mathcal{M}$. This mask can modulate the influence of the consensus vector $\psi_{a}$ on action generation by obscuring it through masked attention. We achieve this through masked attention, setting the consensus mask $\mathcal{M}=0,\psi_{a}=\varnothing$ to prevent the downstream generation of $u_{a}$ from being guided by the consensus representation. Conversely, when $\mathcal{M}=1$, the consensus representation is incorporated alongside the agent’s observation during action generation. During the training phase, the consensus mask $\mathcal{M}$ is sampled from a Bernoulli distribution with a probability $p_{\mathcal{M}=1}$. We set a fixed value $p_{\mathcal{M}=1}=0.2$ during training, resulting in a higher proportion of training samples for consensus-guided over unguided exploration. In execution phase, we adjust $p_{\mathcal{M}=1}=1.0$, encouraging agents to prioritize consensus-driven actions. It’s noteworthy that a recent work [34] demonstrates that consistency policies can achieve policy improvement with precise guidance. In MARL, given the joint policy’s complexity and the environment’s non-stationarity, precise guidance can more significantly enhance the agents’ policy improvement.
@@ IV-C Self-reference Mechanism
Due to the nature of generative models, the consistency policy may generate invalid actions during the initial stages of training [40]. In order to constrain the generated actions, we propose a self-reference mechanism, inspired by self-imitation learning [41]. The self-reference mechanism stores previous experiences in a replay buffer and learns to replicate actions that led to high returns in the past. For each agent, we collect observation-action pairs $(o_{a},u_{a})$, along with their corresponding returns, and store them in the replay buffer $\mathcal{B}$. We then sample entries $(\bm{o},\bm{u},R)$ with empirical returns exceeding the action-value estimates $Q(\bm{o},u_{a})$, creating a reference buffer +=$\mathcal{B}_{+}=\{(\bm{o},\bm{u},R)|R\geq Q(\bm{o},u_{a})\}$, where $u_{a}$ is generated from the current policy $\pi_{\theta_{a}}$. With $\mathcal{B}_{+}$, the consistency policy of agent $a$ can be trained with a reference loss function:
@@
where $\lambda(\cdot)$ denotes a step-dependent weight function, $u^{\tau_{n}}=u+\tau_{n}z$, $z\sim\mathcal{N}(0,I)$ and $d(\cdot,\cdot)$ represents the distance metric. Additionally, $f_{\theta^{\intercal}}$ is the exponential moving average of $f_{\theta}$, introduced for training stability. The self-reference mechanism in our algorithm offers three benefits: a) As generative methods are prone to generating invalid actions [40], the self-reference mechanism acts as a policy constraint, ensuring more reliable action selection. b) Learning from successful transitions can accelerate exploration during the early training stages. c) self-imitation learning can be viewed as a lower bound of soft Q-learning [42], promoting policy improvement. The pseudocode of CPEG is illustrated in Algorithm 1.
@@ 4       Reset the environment and receive initial joint observation $\bm{o}$;
             // Consensus guidance
6             For each agent $a$, infer consensus $\psi_{a}$ by Eq. 8;
7             Sample consensus mask =$\mathcal{M}=\{0,1\}$;
8             Generate $u_{a}$ by $\pi_{\theta_{a}}$ with consensus $\psi_{a}$, consensus mask $\mathcal{M}$, and observation $o_{a}$;
9             Update consensus learner by Eqs. 9, 10 and 11;
@@ 16                   Constrain policy $\pi_{\theta_{a}}$ by self-reference mechanism LABEL:eq:ref;
Algorithm 1 Leveraging Consistency Policy with Consensus Guidance for Multi-agent Exploration
@@ V Experiments
In this section, we empirically evaluate our method to answer the following questions: (1) Does CPEG effectively contribute to exploration and outperform baselines (See Section V-B, Section V-D)? (2) Can consistency policies alleviate the time-consuming issue associated with diffusion policies (See Section V-C)? (3) How does the consensus learner play a role in cooperative exploration (See Section V-D)? (4) Whether consensus guidance and self-reference mechanism contribute collectively to the final performance (See Section V-E)?
@@
Scenarios CPEG (Ours) HATD3 [43] MAPPO [9] HASAC [43] MAT [44] CMAVEN [24]
Dense Reward Navigation -68.1$\pm$2.0 -66.7$\pm$9.1 -84.7$\pm$9.7 -68.6$\pm$8.3 -66.9$\pm$6.8 -88.3$\pm$11.2
Reference -11.9$\pm$1.7 -14.3$\pm$4.2 -32.1$\pm$3.8 -12.6$\pm$2.7 -15.9$\pm$2.9 -37.6$\pm$4.5
Reacher4 (2x1) -22.1$\pm$0.4 -22.3$\pm$0.8 -22.9$\pm$0.5 -21.8$\pm$0.2 -23.0$\pm$0.2 -26.9$\pm$0.3
HalfCheetah (2x3) 6954.9$\pm$466.8 5651.3$\pm$441.3 5463.1$\pm$392.7 7541.3$\pm$472.6 6711.6$\pm$459.8 4526.8$\pm$314.2
Hopper (3x1) 3593.4$\pm$257.1 3349.9±$\pm 187.6$ 3228.3$\pm$194.9 3613.2$\pm$201.4 2745.8$\pm$174.1 2201.4$\pm$222.8
HalfCheetah (6x1) 5692.7$\pm$339.4 4782.4$\pm$438.6 4083.1$\pm$352.1 5535.4$\pm$392.4 5057.4$\pm$334.7 3343.6$\pm$366.2
Average 2689.8$\pm$177.9 2279.9$\pm$180.3 2105.8$\pm$158.9 2764.4$\pm$215.4 2401.4$\pm$195.7 1983.8$\pm$153.2
Sparse Reward Navigation 19.3$\pm$0.3 8.9$\pm$1.3 4.2$\pm$0.8 10.4$\pm$0.6 3.9$\pm$0.9 4.9$\pm$1.8
Reference 18.1$\pm$0.5 10.4$\pm$0.6 3.3$\pm$0.8 10.2$\pm$0.4 3.8$\pm$0.7 8.1$\pm$0.9
Reacher4 (2x1) 50.0$\pm$0.0 36.8$\pm$0.1 24.4$\pm$0.5 38.9$\pm$0.6 37.6$\pm$0.8 28.5$\pm$2.4
HalfCheetah (2x3) 965.0$\pm$8.0 805.0$\pm$12.6 703.7$\pm$27.9 801.2$\pm$10.9 733.7$\pm$14.1 629.1$\pm$13.2
Hopper (3x1) -4.8$\pm$0.7 -33.6$\pm$3.8 -27.9$\pm$2.7 -16.9$\pm$0.9 -20.3$\pm$4.3 -18.4$\pm$3.3
HalfCheetah (6x1) 704.3$\pm$9.8 501.6$\pm$10.2 533.6$\pm$21.9 622.0$\pm$16.4 578.1$\pm$18.2 480.7$\pm$10.3
Average 291.9$\pm$3.8 221.5$\pm$4.7 206.8$\pm$9.0 244.3$\pm$13.9 222.7$\pm$6.5 188.7$\pm$5.3
@@ V-A Experimental Settings
  • Hopper (3x1): This environment features a Hopper in MuJoCo, controlled by agents operating its three primary segments: torso, thigh, and leg. The agents must collaborate effectively to propel the Hopper forward, necessitating intricate coordination as each segment’s movement affects the entire structure’s balance and progress. In sparse-reward settings, the agent receives a reward only upon advancing a specified distance, which places demands on cooperative exploration.
  • HalfCheetah (2x3) and HalfCheetah (6x1): In this environment, the HalfCheetah is a robotic model that resembles a simplified, two-dimensional cheetah. These scenarios involve two and six agents, each representing a front or rear leg. In this shared environment, the agents engage in interactions with the primary objective of advancing forward without falling. These environments parallel the Hopper (3x1) in that agents are rewarded in sparse-reward settings only when they achieve forward movement over a certain distance [45].
@@
Baselines
We compared our results with several baselines as follows. HATD3 and HASAC are strong continuous control multi-agent algorithms, which are newly extended from TD3 and SAC by HARL [43]. MAPPO [9] stands as a popular and effective MARL algorithm in the continuous domain, generating actions from Gaussian distribution. MAT [44] addresses the multi-agent challenge as a sequential modeling issue, using the Transformer [47] to generate complex actions. Notably, we observe that few MARL methods focus on exploration within continuous action spaces. Consequently, we extend the classic MARL exploration algorithm MAVEN [24] into its continuous version, termed CMAVEN. Drawing inspiration from FACMAC [23], CMAVEN adopts the cross-entropy method (CEM) — a sampling-based, derivative-free heuristic search strategy—for approximate greedy action selection. CEM has demonstrated effectiveness in identifying near-optimal solutions in nonconvex Q-networks for single-agent robotic control tasks.
Training Details and Hyper-parameters
We conduct five independent runs with different random seeds in all scenarios. The experiments in this study were conducted using five distinct seeds on a hardware setup comprising 2 NVIDIA RTX A6000 GPUs and 1 AMD EPYC CPU. We utilized official implementations of baseline algorithms, adhering strictly to their original hyper-parameters. The employed consistency policy involves a multi-layer perceptron (MLP), which processes state inputs to generate corresponding actions. Specifically, our policy networks incorporate a 3-layer MLP architecture with a hidden size of 256, using the Mish activation function. The consensus learner consists of a MLP, a consensus codebook module, and another MLP. The MLP network is a 128-dimensional fully-connected layer, followed by a GELU activation, and followed by another 128-dimensional fully-connected layer.
For the consistency policy, we define the diffusion step $k$ within the range of $[0.002,80.0]$, setting the number of sub-intervals $M=40$. Following Karras diffusion model, the sub-interval boundaries are determined with the formula $k_{i}=\left(\epsilon^{\frac{1}{\rho}}+\frac{i-1}{M-1}\left(T^{\frac{1}{\rho}}-\epsilon^{\frac{1}{\rho}}\right)\right)^{\rho}$, where $\rho=7$. The Euler method is employed as the ODE solver. In the consensus learner, we set the number of discrete consensus in codebook $K=5$ in different scenarios, and $\beta=0.2$. The performance of consensus learning is not sensitive to the setting of $K$. Regarding performance evaluation, we pause training every $M$ steps and evaluate for $N$ episodes with action selection. The $(M,N)$ in MPE and MAMuJoCo are $(1000,20)$ and $(10000,40)$, resepcetively. The training objective is optimized by Adam with a learning rate of $5\times 10^{-4}$.
V-B Performance Comparison
We first compare CPEG with baselines across six scenarios. Table I outlines the comparison between our approach and various baseline algorithms in both dense reward and sparse reward settings. In dense-reward environments, CPEG demonstrates performance slightly below or comparable to baselines, surpassing them in specific tasks. In the context of online MARL in POMDPs, the theoretically optimal policy tends to be deterministic, suggesting that expressive models often revert to unimodal to approach optimality. Consequently, expressiveness primarily aids in exploration rather than contributes to the final optimal policy convergence. In sparse-reward environments requiring exploration, CPEG approach exhibits a significant performance advantage, outperforming baseline algorithms by 20%. While methods like HATD3, MAPPO, and MAT show proficiency in dense reward contexts, their limited focus on exploration hinders their effectiveness in reaching sparse reward states. Although the HASAC method improves exploration capability through entropy regularization, the disparity between its employed Gaussian policy and the assumed theoretical Boltzmann policy leads to local optima, thereby constraining exploration. Notably, despite being tailored for MARL exploration, CMAVEN struggles considerably in continuous action spaces, resulting in underwhelming performance. In contrast, our proposed method, employing the consistency policy with consensus guidance and a self-inference mechanism, fosters more efficient exploration and significantly enhances performance.
V-C Time Efficiency
TABLE II: Performance (higher is better) and time efficiency (lower is better) comparison. We compare CPEG, HATD3, and diffusion-based methods in three sparse-reward scenarios, where time denotes the duration of a single iteration.
@@ V-C Time Efficiency
HATD3 CPEG CPEG-DM CPEG-DPM HATD3 CPEG CPEG-DM CPEG-DPM

### v2

@@ IV-B Intention Guidance
The encoder $z_{e}(\cdot;\alpha_{e})$, parameterized by $\alpha_{e}$, encodes the observation of agent $o_{a}$ into an embedding $z_{e,a}\in\mathbb{R}^{m}$, matching the length of the intention representation $m$. In practice, we use historical observations as inputs during each agent’s execution to capture sufficient information about the global state.
Figure 3: The framework of the Intention Learner. The intention learner consists of an observation encoder, an intention codebook, and a state decoder, which learns discrete intention representations by reconstructing states. It operates differently during the training and execution process.
Intention Codebook
The process of inferring intention for each input observation proceeds as follows. First, the embedding $z_{e,a}$ is mapped to a discrete index $k_{a}$ of the codebook $\mathcal{E}$:
@@
where $e_{j}\in\mathcal{E}$ is a codebook vector, and $K$ denotes the size of the codebook. Notably, the performance of intention learning is relatively insensitive to the choice of $K$, allowing us to set the number of discrete intentions in codebook $K=5$ across all scenarios. Based on this discrete index $k_{a}$, the discrete intention representation for the agent $a$ is derived as $\psi_{a}=e_{k_{a}}$. The loss function for the intention inference process is defined as:
@@
where $\mathrm{sg}$ denotes the straight-through gradient operation [43], enabling back-propagation training of $\operatorname*{\arg\!\min}$ operations, and $\beta$ is a hyperparameter balancing the degree of alignment between the codebook and input embeddings. In practice, CPIG uses the exponential moving average (EMA) to update the codebook, instead of directly learning them as parameters. Specifically, EMA updates the codebook $e_{k_{a}}$ by replacing it with a weighted combination of its previous value $e_{k_{a}}$ and the corresponding vector $z_{e,a}$:
@@
To develop the ability to deduce intention from local observations, we aggregate the intention $\psi_{a}$ from all agents together and process them through a state decoder $z_{d}(\cdot;\alpha_{d})$, parameterized by $\alpha_{d}$, to reconstruct the global state $\hat{s}$. We train the observation encoder and state decoder together by maximizing the log-likelihood of the reconstruction:
@@
which encourages the encoding and decoding processes of the intention learner to capture the key information from partial observations for accurate intention deduction. In addition, the reconstruction loss gradient is also passed to the encoder for training $\alpha_{e}$ by straight-through gradient estimation.
Mask Intention Exploration
In this work, we employ two distinct exploration strategies: unguided exploration and intention-guided exploration, aiming to balance exploration and exploitation to some degree. Unguided exploration allows agents to take action according to their current policies without specific intention guidance, emphasizing their self-directed exploration capabilities. However, its effectiveness is limited by the initial policy quality and the inherent complexities of MAS, neglecting potential cooperative advantages. Considering this, we introduce another exploration strategy, intention-guided exploration designed to steer the cooperation efforts of agents. The intention-guided exploration instructs agents to take actions based on the consistency policy guided by their intention. Through the guidance of shared intention among different agents, consistency policies can generate cooperative exploration behaviors. Meanwhile, to incorporate intention guidance $\psi$, the consistency function $F_{\theta}(o,u,\tau)$ in Eq. 4 is modified to $F_{\theta}(o,u,\psi,\tau)$.
To implement the above concept and combine the two exploration mechanisms, we introduce a binary intention mask $\mathcal{M}$. This mask can modulate the influence of the intention vector $\psi_{a}$ on action generation by obscuring it through masked attention. We achieve this through masked attention, setting the intention mask $\mathcal{M}=0,\psi_{a}=\varnothing$ to prevent the downstream generation of $u_{a}$ from being guided by the intention representation. Conversely, when $\mathcal{M}=1$, the intention representation is incorporated alongside the agent’s observation during action generation. During the training phase, the intention mask $\mathcal{M}$ is sampled from a Bernoulli distribution with a probability $p_{\mathcal{M}=1}$. We set a fixed value $p_{\mathcal{M}=1}=0.2$ during training, resulting in a higher proportion of training samples for intention-guided over unguided exploration. In execution phase, we adjust $p_{\mathcal{M}=1}=1.0$, encouraging agents to prioritize intention-driven actions. It’s noteworthy that a recent work [37] demonstrates that consistency policies can achieve policy improvement with precise guidance. In MARL, given the joint policy’s complexity and the environment’s non-stationarity, precise guidance can more significantly enhance the agents’ policy improvement.
@@ IV-C Self-reference Mechanism
Due to the nature of generative models, the consistency policy may generate invalid actions during the initial stages of training [44]. In order to constrain the generated actions, we propose a self-reference mechanism, inspired by self-imitation learning [45]. The self-reference mechanism stores previous experiences in a replay buffer and learns to replicate actions that led to high returns in the past. For each agent, we collect observation-action pairs $(o_{a},u_{a})$, along with their empirical return $R$, and store them in the replay buffer $\mathcal{B}$. We then sample entries $(\bm{o},\bm{u},R)$ with empirical returns exceeding the action-value estimates $Q(\bm{o},u_{a})$, creating a reference buffer +=$\mathcal{B}_{+}=\{(\bm{o},\bm{u},R)|R\geq Q(\bm{o},u_{a})\}$, where $u_{a}$ is generated from the current policy $\pi_{\theta_{a}}$. By uniformly sampling $(o_{a},u_{a})$ from $\mathcal{B}_{+}$, the consistency policy of agent $a$ can be trained using a reference loss function:
@@
where $\lambda(\cdot)$ denotes a step-dependent weight function, $u^{\tau_{n}}=u+\tau_{n}z$ and $z\sim\mathcal{N}(0,I)$. $d(\cdot,\cdot)$ represents the distance metric and we use $l_{2}$ distance $d(x,y)=\|x-y\|^{2}_{2}$. Additionally, $f_{\theta^{\intercal}}$ is the exponential moving average of $f_{\theta}$, introduced for training stability. The self-reference mechanism in our algorithm offers three benefits: a) As generative methods are prone to generating invalid actions [44], the self-reference mechanism acts as a policy constraint, ensuring more reliable action selection. b) Learning from successful transitions can accelerate exploration during the early training stages. c) self-imitation learning can be viewed as a lower bound of soft Q-learning [46], promoting policy improvement. The pseudocode of CPIG is illustrated in Algorithm 1.
@@ 4       Reset the environment and receive initial joint observation $\bm{o}$;
             // Intention guidance
6             For each agent $a$, infer intention $\psi_{a}$ by Eq. 8;
7             Sample intention mask =$\mathcal{M}=\{0,1\}$;
8             Generate $u_{a}$ by $\pi_{\theta_{a}}$ with intention $\psi_{a}$, intention mask $\mathcal{M}$, and observation $o_{a}$;
9             Update intention learner by Eqs. 9, 10 and 11;
@@ 16                   Constrain policy $\pi_{\theta_{a}}$ by self-reference mechanism LABEL:eq:ref;
Algorithm 1 Leveraging Consistency Policy with Intention Guidance for Multi-agent Exploration
@@ V Experiments
In this section, we empirically evaluate our method to answer the following questions: (1) Does CPIG effectively contribute to exploration and outperform baselines (See Section V-B, Section V-D)? (2) Can consistency policies alleviate the time-consuming issue associated with diffusion policies (See Section V-C)? (3) How does the intention learner play a role in cooperative exploration (See Section V-D)? (4) Whether intention guidance and the self-reference mechanism contribute collectively to the final performance (See Section V-E)?
@@
Scenarios CPIG (Ours) HATD3 [47] MAPPO [10] HASAC [47] MAT [48] CMAVEN [27] CMAE [20]
Dense Reward Navigation -68.1$\pm$2.0 -66.7$\pm$9.1 -84.7$\pm$9.7 -68.6$\pm$8.3 -66.9$\pm$6.8 -88.3$\pm$11.2 -82.4$\pm$10.9
Reference -11.9$\pm$1.7 -14.3$\pm$4.2 -32.1$\pm$3.8 -12.6$\pm$2.7 -15.9$\pm$2.9 -37.6$\pm$4.5 -21.1$\pm$3.9
Reacher4 (2x1) -22.1$\pm$0.4 -22.3$\pm$0.8 -22.9$\pm$0.5 -21.8$\pm$0.2 -23.0$\pm$0.2 -26.9$\pm$0.3 23.4$\pm$0.6
HalfCheetah (2x3) 6954.9$\pm$466.8 5651.3$\pm$441.3 5463.1$\pm$392.7 7541.3$\pm$472.6 6711.6$\pm$459.8 4526.8$\pm$314.2 5783.6$\pm$443.6
Hopper (3x1) 3593.4$\pm$257.1 3349.9±$\pm 187.6$ 3228.3$\pm$194.9 3613.2$\pm$201.4 2745.8$\pm$174.1 2201.4$\pm$222.8 2328.8$\pm$146.2
HalfCheetah (6x1) 5692.7$\pm$339.4 4782.4$\pm$438.6 4083.1$\pm$352.1 5535.4$\pm$392.4 5057.4$\pm$334.7 3343.6$\pm$366.2 3253.0$\pm$401.7
Average 2689.8$\pm$177.9 2279.9$\pm$180.3 2105.8$\pm$158.9 2764.4$\pm$215.4 2401.4$\pm$195.7 1653.1$\pm$153.2 1880.9$\pm$167.8
Sparse Reward Navigation 19.3$\pm$0.3 8.9$\pm$1.3 4.2$\pm$0.8 10.4$\pm$0.6 3.9$\pm$0.9 4.9$\pm$1.8 7.6$\pm$1.1
Reference 18.1$\pm$0.5 10.4$\pm$0.6 3.3$\pm$0.8 10.2$\pm$0.4 3.8$\pm$0.7 8.1$\pm$0.9 9.3$\pm$0.8
Reacher4 (2x1) 50.0$\pm$0.0 36.8$\pm$0.1 24.4$\pm$0.5 38.9$\pm$0.6 37.6$\pm$0.8 28.5$\pm$2.4 36.1$\pm$1.0
HalfCheetah (2x3) 965.0$\pm$8.0 805.0$\pm$12.6 703.7$\pm$27.9 801.2$\pm$10.9 733.7$\pm$14.1 629.1$\pm$13.2 742.7$\pm$16.8
Hopper (3x1) -4.8$\pm$0.7 -33.6$\pm$3.8 -27.9$\pm$2.7 -16.9$\pm$0.9 -20.3$\pm$4.3 -18.4$\pm$3.3 -21.9$\pm$2.3
HalfCheetah (6x1) 704.3$\pm$9.8 501.6$\pm$10.2 533.6$\pm$21.9 622.0$\pm$16.4 578.1$\pm$18.2 480.7$\pm$10.3 541.5$\pm$17.2
Average 291.9$\pm$3.8 221.5$\pm$4.7 206.8$\pm$9.0 244.3$\pm$13.9 222.7$\pm$6.5 188.7$\pm$5.3 219.3$\pm$ 6.5
@@ V-A Experimental Settings
  • Hopper (3x1): This environment features a Hopper in MuJoCo, controlled by agents operating its three primary segments: torso, thigh, and leg. Effective forward propulsion requires precise coordination among agents, as the movement of each segment affects the balance and overall progress of the Hopper. In sparse-reward settings, agents receive rewards only after the Hopper advances a specified distance, making cooperative exploration essential.
  • HalfCheetah (2x3) and HalfCheetah (6x1): In this environment, the HalfCheetah is a robotic model that resembles a simplified, two-dimensional cheetah. The two scenarios feature either two agents, each controlling a front or rear leg, or six agents, each controlling a single joint. In both cases, the agents must coordinate their actions to achieve forward movement while avoiding falls. Similar to the Hopper (3x1) environment, rewards in sparse-reward settings are granted only when the agent successfully moves forward over a specified distance [49].
@@
TABLE II: Performance (higher is better) and time efficiency (lower is better) comparison. We compare CPIG, HATD3, and diffusion-based methods in three sparse-reward scenarios, where time denotes the duration of a single iteration.
@@
HATD3 CPIG CPIG-DM CPIG-DPM HATD3 CPIG CPIG-DM CPIG-DPM
@@
Baselines
We compared our results with several baselines as follows. HATD3 and HASAC are strong continuous control multi-agent algorithms, which are newly extended from TD3 and SAC by HARL [47]. MAPPO [10] stands as a popular and effective MARL algorithm in the continuous domain, generating actions from Gaussian distribution. MAT [48] addresses the multi-agent challenge as a sequential modeling issue, using the Transformer [51] to generate complex actions. Notably, we observe that few MARL methods focus on exploration within continuous action spaces. Consequently, we extend the classic MARL exploration algorithm MAVEN [27] into its continuous version, termed CMAVEN. Drawing inspiration from FACMAC [26], CMAVEN adopts the cross-entropy method (CEM) — a sampling-based, derivative-free heuristic search strategy—for approximate greedy action selection. CEM has demonstrated effectiveness in identifying near-optimal solutions in nonconvex Q-networks for single-agent robotic control tasks. In addition, we introduce CMAE [20] and adopt the settings described in the original paper, utilizing hash-based counting techniques [52] to extend the algorithm to continuous state spaces.
Training Details and Hyper-parameters
We conduct five independent runs with different random seeds in all scenarios. The experiments in this study were conducted using five distinct seeds on a hardware setup comprising 2 NVIDIA RTX A6000 GPUs and 1 AMD EPYC CPU. We utilized official implementations of baseline algorithms, adhering strictly to their original hyper-parameters. The employed consistency policy involves a multi-layer perceptron (MLP), which processes state inputs to generate corresponding actions. Specifically, our policy networks incorporate a 3-layer MLP architecture with a hidden size of 256, using the Mish activation function. The intention learner consists of a MLP, an intention codebook module, and another MLP. The MLP network is a 128-dimensional fully-connected layer, followed by a GELU activation, and followed by another 128-dimensional fully-connected layer.
For the consistency policy, we define the diffusion step $k$ within the range of $[0.002,80.0]$, setting the number of sub-intervals $M=40$. Following Karras diffusion model, the sub-interval boundaries are determined with the formula $k_{i}=\left(\epsilon^{\frac{1}{\rho}}+\frac{i-1}{M-1}\left(T^{\frac{1}{\rho}}-\epsilon^{\frac{1}{\rho}}\right)\right)^{\rho}$, where $\rho=7$. The Euler method is employed as the ODE solver. In the intention learner, we set the number of discrete intentions in codebook $K=5$ in different scenarios, and $\beta=0.2$. The performance of intention learning is not sensitive to the setting of $K$. Regarding performance evaluation, we pause training every $M$ steps and evaluate for $N$ episodes with action selection. The $(M,N)$ in MPE and MAMuJoCo are $(1000,20)$ and $(10000,40)$, resepcetively. The training objective is optimized by Adam with a learning rate of $5\times 10^{-4}$.
V-B Performance Comparison
We first compare CPIG with baselines across six scenarios. Table I outlines the comparison between our approach and various baseline algorithms in both dense reward and sparse reward settings. In dense-reward environments, CPIG demonstrates performance slightly below or comparable to baselines, surpassing them in specific tasks. In the context of online MARL in POMDPs, the theoretically optimal policy tends to be deterministic, suggesting that expressive models often revert to unimodal to approach optimality. Consequently, expressiveness primarily aids in exploration rather than contributes to the final optimal policy convergence. In sparse-reward environments requiring exploration, CPIG approach exhibits a significant performance advantage, outperforming baseline algorithms by 20%. While methods like HATD3, MAPPO, and MAT show proficiency in dense reward contexts, their limited focus on exploration hinders their effectiveness in reaching sparse reward states. Although the HASAC method improves exploration capability through entropy regularization, the disparity between its employed Gaussian policy and the assumed theoretical Boltzmann policy leads to local optima, thereby constraining exploration. Notably, despite being tailored for MARL exploration, CMAVEN and CMAE struggle considerably in continuous action or state spaces, resulting in underwhelming performance. In contrast, our proposed method, employing the consistency policy with intention guidance and a self-inference mechanism, fosters more efficient exploration and significantly enhances performance.
V-C Time Efficiency